OpenStudy (anonymous):

Suppose that in one week, your company takes in $2,220 in sales at one of its cell phone stores. The price breakdown per phone is as follows: Cell phone model A4: $35 Smartphone Z20: $50 Suppose that a total of 51 phones were sold. this is the supposed problem with the method of solving in a graph

OpenStudy (anonymous):

I would start this part of the assignment by identifying the variables. I would set it up as follows: x = Cell phone model A4 y = Smartphone z20 Now I must set up a system of equations to figure out the solution. x + y = 51 $35x + $50y = $2220 I would solve this using the substitution method. Since more of the z20’s were purchased than the A4’s, we will now substitute y in either equation to solve for x. x + (x + 7) = 51 find x. -7 = -7 2x = 44 now, 44 divided by 2 gives us x = 22. Now we plug the value of x into the other equation to find the value of y. This is known as distribution. $35(22) + $50y = $2,220 Now time to simplify. $35 * 22 = $770 + 50y = $2,220 Now we got to get the y term by itself so: to solve a linear equation I would now subtract the $770 from both sides of the equation. The rule with linear equations is what we do to one side of the equal sign you do to the other. So: $770 + 50y = $2,220 -$770 -$770 50y = $1450 Now since y is multiplied, we need to divide $1450 by 50 to find the value of y. $1450 ÷ 50 = 29 so y = 29 So in summary if x = cell phone model A4 and y = smartphone z20’s… We have sold: 22 of the cell phone model A4 and 29 smartphone z20’s based on taking in $2,220 in sales for the day.

OpenStudy (anonymous):

thanks u helped so much i appreciate it :)